# Rigorous and Generalized Proof of Security of Bitcoin Protocol with Bounded Network Delay

### v1

@@ IV-B The punctured arrival process
Figure 2: Figure representing the punctured arrival process. $B$ is the time-length of the part of each interval that is not deleted. Without loss of generality, we start this process at a time $t=0$. Recall that $\Delta$ is the maximum delay of the $\Delta$-delay model, and also is the length of each puncture. In this process, all honest blocks that arrive during these length $\Delta$ intervals are deleted. $B(i)$ is the $i$th total punctured score, which is the total score of the fully-delayed chain after the honest blocks in the punctured intervals are deleted. $S_{b,i}$ is the $i$th punctured interval score, which is the score growth of the fully-delayed honest blocks in each length $B$ interval. Observe that, due to the length $\Delta$ deletion, each $S_{B,i}$ is independent and identically distributed.
@@
Let $B(n)=S_{B}(nB+n\Delta)$, which is a random variable that represents the score of the punctured chain at the end of the $n$th puncture, which we call the $n$th total punctured score.
@@
$B(n)=\sum_{i=1}^{n}S_{B,i}.$
Using the above equation, and substituting the definition of $B(n)$ into (5) gives us, for each integer $n>0$:
$S(Bn+\Delta n)\geq B(n)=\sum_{i=1}^{n}S_{B,i}$
@@
$P\left(\left[\frac{B(n)}{n}\geq B\lambda_{h}-B\epsilon\right]\text{ for all $n>0$}\right)$
@@
Substituting the definition of $B(n)$ on the left and summing up $n$ constants on the right gives us:
$B(n)>Bn\lambda_{h}-nB\epsilon\text{ for all integers $n>0$}$
@@ A-A Counterexample
Appendix B The Chernoff Bound Argument
@@
Suppose the blockrate for honest non-delayed blocks for each block type $i$ is $h_{i}$, and the score of each block type is $c_{i}$. Hence, the non-delayed chain has score growth rate $\lambda^{\prime}=\sum c_{i}h_{i}$. We let the fully-delayed growth rate of blocks of type $i$ be $h_{d}^{i}$ (note that, by the same argument as Lemma 16, this growth rate exists). Hence, the fully-delayed honest score growth rate is given by $\lambda_{h}=\sum c_{i}h_{d}^{i}$.
@@
Probability that score growth of one block type is close to or below below average scales as $A\exp(-ct^{\prime}$)
We consider the probability that, in the fully delayed chain, blocks of a particular type have number of arrivals close to or below average. We show that this probability goes to 0 exponentially in the time window $t$.
Without loss of generality, let us consider blocks of type $1$. Let $Y_{d}$ be the random variable representing the number of blocks of this type that arrive in the fully delayed chain in a length of time $t^{\prime}$. Let $\mu_{d}$ be the average number of arrivals of blocks of this type in the fully-delayed chain in this interval. Let $h_{d}$ the fully-delayed blockrate for type $1$ blocks (in blocks per second) in the full chain, (where in this subsection we suppress the indices $i$ for notational clarity).
Note 1: In any particular finite interval of length $t^{\prime}$, the expected number of blocks of this type is close to but may not be exactly $h_{d}t^{\prime}$ (as this can depend on, for example, blocks arriving within $\Delta$ of the end of the interval). However, for the same reasoning as Lemma 22, because $h_{d}$ is the limit of these average arrival rates as $t^{\prime}$ approaches infinity, then for any $\delta>0$, and for sufficiently large $t$’, $(1-\delta)h_{d}t^{\prime}\leq\mu_{d}$. We shall use this observation later in our proof.
Directly applying the Chernoff bound formula for $Y_{d}$ gives us for all $y<\mu_{d}$ and $s<0$:
$P(Y_{d}\leq y)\leq\frac{E[e^{sY_{d}}]}{e^{sy}}$
(as a reference, see [8]).
We do not have a formula for the moment generating function ($E[e^{sY_{d}}]$) of $Y_{d}$. However, let $Y$ be the number of arrivals of type $1$ blocks without delays in this interval. This process has average number of arrivals $\mu=t^{\prime}h_{i}$. Obviously, since no blocks get erased in the non-delayed process:
$Y_{d}$ $\leq Y$
$sY_{d}$ $\leq sY.$
Hence:
$e^{sY_{d}}\leq e^{sY}$
and thus
$E\left(e^{sY_{d}}\right)\leq E\left(e^{sY}\right).$
Substituting this bound into the formula above, and using the well-known formula for the $E\left(e^{sY}\right)$ [8]:
$P(Y_{d}\leq y)\leq\frac{e^{\mu(e^{s}-1)}}{e^{sy}}.$
Now we simply let $s=\ln(y/\mu_{d})<0$ into the formula. This gives us:
$P(Y_{d}\leq y)\leq\frac{\exp(\mu\left(\exp\left(\ln(y/\mu_{d})\right)-1\right))}{\exp\left(\ln(y/\mu_{d})y\right)}=\frac{\exp(\mu\left(\frac{y}{\mu_{d}}-1\right))}{\exp\left(\ln(y/\mu_{d})y\right)}$
Now let let $y=(1-\delta)h_{d}t^{\prime}$ (which for all $\delta$ and sufficiently large $t^{\prime}$ is less than $\mu_{d}$, and so is a valid choice for $y$):
$P\left(Y_{d}\leq(1-\delta)h_{d}t^{\prime}\right)=\frac{\exp(\mu\left(\frac{(1-\delta)h_{d}t^{\prime}}{\mu_{d}}-1\right))}{\exp\left(\ln(\frac{(1-\delta)h_{d}t^{\prime}}{\mu_{d}})\left((1-\delta)h_{d}t^{\prime}\right)\right)}$
which by inspection we see scales as $A\exp(-ct^{\prime}\ln t)\leq A\exp(-ct^{\prime}$).
Hence:
$P\left(Y_{d}\leq(1-\delta)h_{d}t^{\prime}\right)\leq A\exp(-ct^{\prime}).$
B-A Union Bound Argument for Adversary Chain Dominating Event
Above, we showed that the probability that a particular block type has below average rate of arrivals in our length $t^{\prime}$ interval scales as $A\exp(-ct^{\prime})$.
A basic Chernoff bound argument can also show that the probability that the adversary has arrivals less than or equal to $((1+\delta)\lambda_{a}t^{\prime}$ scales as $A\exp(-ct^{\prime})$.
Recall we are in a region in which $\lambda_{h}>\lambda_{a}$. Hence,
@@ B-A Union Bound Argument for Adversary Chain Dominating Event
$(1-\delta)t^{\prime}\lambda_{h}>\left(1+\delta\right)t^{\prime}\lambda_{a}$ (16)
for sufficiently small $\delta$. Let us consider such a $\delta$ below.
Observe that if all honest block types have $(1-\delta)h_{d}^{i}t^{\prime}$ or greater arrivals in this interval, and the adversary has less than $(1+\delta)\lambda_{a}t^{\prime}$, then:
honest chain score growth in interval $\geq\sum_{i=1}^{n}(1-\delta)c_{i}h_{d}^{i}t^{\prime}=$
$=$ $(1-\delta)t^{\prime}\sum_{i=1}^{n}c_{i}h_{d}^{i}$
$=$ $(1-\delta)t^{\prime}\lambda_{h}$
$>$ $\left(1+\delta\right)t^{\prime}\lambda_{a}$
where in the last line we use 16. But we assumed adversary arrivals are less than $\left(1+\delta\right)t^{\prime}\lambda_{a}$.
Hence, in order for the adversary chain to dominate in this case, at least one of the block types must have less than average growth rate, or the adversary has more than average growth rate. Each of these occur with probability that scales at least as $A\exp(-ct^{\prime})$. By union bound, the probability of such a dominating event occuring also scales as $A\exp(-ct^{\prime})$.
@@
B-B The Integration Argument: The Probability any Block in Interval is Dominated by Long Adversary Chain

### v2

@@ IV-B The punctured arrival process
Figure 2: Figure representing the punctured arrival process. $B$ is the time-length of the part of each interval that is not deleted. Without loss of generality, we start this process at a time $t=0$. Recall that $\Delta$ is the maximum delay of the $\Delta$-delay model, and also is the length of each puncture. In this process, all honest blocks that arrive during these length $\Delta$ intervals are deleted. $B_{t}(i)$ is the $i$th total punctured score, which is the total score of the fully-delayed chain after the honest blocks in the punctured intervals are deleted. $S_{b,i}$ is the $i$th punctured interval score, which is the score growth of the fully-delayed honest blocks in each length $B$ interval. Observe that, due to the length $\Delta$ deletion, each $S_{B,i}$ is independent and identically distributed.
@@
Let $B_{t}(n)=S_{B}(nB+n\Delta)$, which is a random variable that represents the score of the punctured chain at the end of the $n$th puncture, which we call the $n$th total punctured score.
@@
$B_{t}(n)=\sum_{i=1}^{n}S_{B,i}.$
Using the above equation, and substituting the definition of $B_{t}(n)$ into (5) gives us, for each integer $n>0$:
$S(Bn+\Delta n)\geq B_{t}(n)=\sum_{i=1}^{n}S_{B,i}$
@@
$P\left(\left[\frac{B_{t}(n)}{n}\geq B\lambda_{h}-B\epsilon\right]\text{ for all $n>0$}\right)$
@@
Substituting the definition of $B_{t}(n)$ on the left and summing up $n$ constants on the right gives us:
$B_{t}(n)>Bn\lambda_{h}-nB\epsilon\text{ for all integers $n>0$}$
@@ A-A Counterexample
Appendix B Probability that Honest Growth Rate Deviates from Average Decays Exponentially
@@
B-A Probability honest growth rate deviates from average is exponential in time $t$
Consider the punctured process defined in Section IV-B.
For any choice of $\epsilon>0$, from Lemma 16 we know that as $B$ gets large its expected value approaches $\lambda_{h}$. So choose $B$ sufficiently large so that
$E(B)=\lambda_{h}-\delta\epsilon$
and
$B(1-\delta\epsilon)+\Delta(\lambda_{h}-\epsilon)<0.$ (16)
The latter inequality can be true so long as $\delta$ is sufficiently small (which occurs as $B$ gets larger) and $B$ sufficiently big.
Let $B_{t}(N)$ be the sum of the scores of the first $n$ punctured intervals.
We have the time at the end of the $n$th puncture is:
$t=(B+\Delta)n$
We let
$n(t)=\left\lfloor\frac{t}{B+\Delta}\right\rfloor$
We additionally let $S_{p}(t)$ be the score of the punctured chain at time $t$.
By Lemma 5, deletion can only decrease the score. Also, truncating to the nearest mutliple of $t$ can also only decrease the score.
$B_{t}(n(t))=B_{t}\left(\left\lfloor\frac{t}{b+\Delta}\right\rfloor\right)\leq S_{p}(t)\leq S(t).$
Consider the event $S(t)\leq t(\lambda_{h}-\epsilon)$, (i.e. that the fully-delayed honest chain at time $t$ has average score growth rate a value $\epsilon>0$ below average.
If we weaken the honest score growth rate by puncturing it and also truncating it at the endpoints of each of the punctured segments, this can only increase the probability that the score growth is below a certain value. Hence,
$P\left[S(t)\leq t(\lambda_{h}-\epsilon)\right]\leq P\left[B_{t}(n)\leq t(\lambda_{h}-\epsilon)\right]$
Note that $\left(n+1\right)(B+\Delta)\geq t$ (where implicitly $n$ is a function of $t$). Observe that replacing $t(\lambda_{h}-\epsilon)$ with a bigger value above can only increase the probability that $B(n)$ is less than it. Hence, replacing $t$ with $\left(n+1\right)(B+\Delta)$ gives us,
$P\left[S(t)\leq t(\lambda_{h}-\epsilon)\right]\leq P\left[B_{t}(n)\leq t(\lambda_{h}-\epsilon)\right]\leq P\left[B_{t}(n)\leq\left(n+1\right)(B+\Delta)(\lambda_{h}-\epsilon)\right]$ (17)
@@ B-A Probability honest growth rate deviates from average is exponential in time $t$
For now we consider understanding the event $\left\{B_{t}(n)\leq(n+1)(B+\Delta)(\lambda_{h}-\epsilon)\right\}$ .
Then, the event above can be expressed as:
$\left\{B_{t}(n)\leq\left(n+1\right)(B+\Delta)(\lambda_{h}-\delta\epsilon-(1-\delta)\epsilon)\right\}$
Expanding the right side and simplifying:
$B_{t}(n)\leq nB(\lambda_{h}-\delta\epsilon)-nB(1-\delta)\epsilon+(n+1)\Delta(\lambda_{h}-\epsilon)$
Moving some terms to the other side of the inequality:
$B_{t}(n)-nB(\lambda_{h}-\delta\epsilon)\leq-nB(1-\delta\epsilon)+(n+1)\Delta(\lambda_{h}-\epsilon)$
Dividing both sides by $n$:
$\frac{B_{t}(n)-n(B)(\lambda_{h}-\delta\epsilon)}{n}\leq-B(1-\delta\epsilon)+\frac{(n+1)\Delta(\lambda_{h}-\epsilon)}{n}.$
Observe that for the right side of the inequality,
$B(1-\delta\epsilon)+\frac{(n+1)\Delta(\lambda_{h}-\epsilon)}{n}\rightarrow-B(1-\delta\epsilon)+\Delta(\lambda_{h}-\epsilon)$
as $n$ gets large. Thus, from 16, the right side is negative for sufficiently large $n$ . Let k-k equal the right side of this inequality. Moreover, recognize that $E(B)=(B)(\lambda_{h}-\delta\epsilon)$. Thus, we are this interested in the event that:
$\frac{B_{t}(n)-n(B)(\lambda_{h}-\delta\epsilon)}{n}<-k.$
so that we are considering:
$P\left[\frac{B_{t}(n)-\mu_{p}}{n}\leq-k\right].$
The Hoeffding inequality states that for random variables $X(n)$ that are independent and bounded in the range $[a,b]$ for some finite $a$ and $b$, and all $t>0$:
$P\left[\frac{\sum_{i=1}^{n}X(i)-E(X(i))}{n}\leq-t\right]\leq\exp\left(-\frac{2nt^{2}}{(b-a)^{2}}\right).$
Observe that the value of $B$ is at least 0 almost surely (corresponding to the event of no honest arrivals at all in the interval), and at most is $b:=\max_{i}(c_{i})\frac{B}{\Delta}$, corresponding to the case where the highest score block type arrives exactly at $\Delta$ after each prior block arrival. Hence, the Hoeffding inequality applies for $n$ sufficiently large, and thus
$P\left[\frac{B_{t}(n)-\mu_{p}}{n}\leq-k\right]\leq\exp\left(-\frac{2nk^{2}}{(b)}\right).$
Substituting this result back into the chain of inequalities in 17, and substituting $n=\left\lfloor\frac{t}{B+\Delta}\right\rfloor$ gives us:
$P\left[S(t)\leq t(\lambda_{h}-\epsilon)\right]\leq A\exp(-ct)$
for some constants $A,c>0$.
B-B This implies probability of a dominating event of length $t$ is exponential
A basic chernoff bound argument can show that the probability that an adversary has average growth rate above average $\lambda_{a}$ decays exponentialky to 0 in length $t^{\prime}$. Since this is just a straightforward application of a Chernoff bound, we omit the proof here.
Recall we are in a region in which $\lambda_{h}>\lambda_{a}$. Hence, for any $t^{\prime}>0$:
$\lambda_{h}-\epsilon>\lambda_{a}+\epsilon$ (18)
for sufficiently small $\epsilon$. It is now easy to see that in order for an adversary to dominate at a time $t^{\prime}$ from an honest block arrival, it must be that either the honest blocks must grow at a rate above $\lambda_{h}-\epsilon$ or the adversary must grow at a rate above $\lambda_{a}+\epsilon$. Each of these occur with exponentially decaying probability, and thus by union bound this event also decays with exponential probability.
@@ B-B This implies probability of a dominating event of length $t$ is exponential
B-C The Integration Argument: The Probability any Block in Interval is Dominated by Long Adversary Chain